# Gas Cost Analysis of Proxy and Diamond Patterns: Towards Trusted Smart Contract Engineering in EVM Blockchains

### v1

@@ Abstract
Blockchain applications are witnessing rapid evolution, necessitating the integration of upgradeable smart contracts. Software patterns have been proposed to summarize upgradeable smart contract best practices. However, research is missing on the comparison of these upgradeable smart contract patterns, especially regarding gas costs related to deployment and execution. This study aims to provide an in-depth analysis of gas costs associated with two prevalent upgradeable smart contract patterns: the Proxy and diamond patterns. The Proxy pattern utilizes a Proxy pointing to a logic contract, while the diamond pattern enables a Proxy to point to multiple logic contracts. We conduct a comparative analysis of gas costs for both patterns in contrast to a traditional non-upgradeable smart contract. We derive from this analysis a theoretical contribution in the form of two consolidated blockchain patterns and a corresponding decision model. By so doing we hope to contribute to the broader understanding of upgradeable smart contract patterns.
@@ 1 Introduction
In the world of blockchain technology, smart contracts stand as keystones to orchestrate reliably and securely digital transactions [20]. Smart contracts are particularly important on Ethereum and similar blockchain platforms, where they are used in diverse areas including digital finance, identity management, or industrial traceability.
@@
As blockchain applications evolve, the imperative for smart contracts to be not only secure but also upgradeable becomes apparent [10]. In this context, a set of upgradeable smart contract patterns, including the Proxy, and Diamond, have surfaced to answer the lack of adaptability of classical smart contracts whose logic cannot be changed once deployed [23, 27]. The Proxy pattern, wielding a proxy contract to delegate calls to an implementation contract, offers a flexible and upgradeable alternative. The Diamond pattern, as proposed in EIP-2535, addresses concerns like maximum contract size, ensuring enhanced maintainability and versioning through multiple implementation contracts.
@@
Research has focused on identifying upgradeable smart contracts pattern families, pointing towards Proxy and Diamond-based strategies [22, 16, 23, 27]. However, the papers do not provide an in-depth analysis of the functional and non-functional properties of these patterns, nor gas costs behaviors. Hence, a research gap is identified regarding a thorough study of the Proxy and Diamond patterns especially in terms of a gas costs analysis.
@@
To address this issue, this paper aims to answer the following research questions: (RQ1) How do Proxy and Diamond patterns differ in terms of gas consumption, scalability, and ease of use? and (RQ2) What implications do these differences hold for developing blockchain applications?
@@
In this paper, we contribute to the literature through a unified approach to compare gas costs in upgradeable smart contracts. We leverage this methodology to provide a comparative gas cost analysis of the Proxy and Diamond patterns in EVM blockchains, compared against a monolithic non-upgradeable smart contract, which is used as baseline. Based on these results, we derive a theoretical contribution in the form of two smart contract patterns adhering to the Alexandrian form format following the standard proposed by Christopher Alexander [2]. These patterns include the results of our comparative analysis and contribute to the broader understanding of upgradeable smart contract patterns and their use in blockchain applications. Based on these patterns, we also propose a decision model for using these patterns that highlights functional and non-functional properties involved in each decision design path.
@@
The remainder of this paper is structured as follows. Section 2 introduces key concepts related to smart contracts and blockchain patterns, and section 3 presents studies already made on linked concepts. Section 4 presents the method used to carry our comparative analysis on Proxy versus Diamond gas costs. Section 5 presents the results of the tests made on the different patterns. Section 6 leverages these findings to propose two consolidated Proxy and Diamond patterns as well as a decision model for using these patterns. Section 7 finally concludes the paper with a summary and a discussion of the results and some considerations for future work.
@@ 2 Background
This section introduces key concepts that will be used in the sequel of this paper revolving around smart contracts and related concepts such as upgradability patterns, EVM, gas and storage.
@@ 2.1 Smart Contracts
A smart contract is a program hosted on a blockchain network [30]. When a smart contract executes, its updated state (or storage) is registered into a transaction that will be stored in the blockchain ledger making it immutable and tamper-proof [5]. More precisely, the final validated results are stored in a Merkle Patricia trie whose nodes correspond to an account or a smart contract. A smart contract comprises both variables and functions. Smart contract functions can execute arbitrary code, access the state of the variables and optionally update them. The default function’s mutability gives right to read and modify state variables. However, function mutability can be constrained to add more security when it comes to accessing these variables. On the one hand, a pure function cannot read nor modify the state of the contract; it is only used to compute a value, often using the parameters passed to it. On the other hand, a view function can read the state of the contract, but cannot modify it.
Smart contracts are immutable, meaning that once deployed, they cannot be modified [31, 9]. This is a security feature, as it prevents malicious actors from modifying the code, so that the contract can be trustable and unbreakable. However, this also means that if a bug is found in the code, it cannot be fixed, and the only solution is to deploy a new contract. Also, if a new feature is added to the contract, the only way to do it is to deploy a new contract. This is a problem for users, as they have to migrate to the new contract, and for developers, as they have to maintain multiple contracts. To address this issue, several patterns have emerged [10, 23], which are discussed later in the paper.
@@ 2.2 Gas and storage in the Ethereum Virtual Machine (EVM)
The EVM is a Turing-complete virtual machine that runs on every node of the Ethereum network and other EVM-based blockchains. It provides a secure and isolated environment for smart contracts execution [18]. Gas is the unit of computation in the EVM, and every operation performed by the EVM has a gas cost associated with it [17]. The user pays this cost in the form of gas fees, to the miner/validator who executes the smart contract. It is the miner’s incentive to execute the smart contract and record the results on the blockchain, but also a way to prevent spam and denial of service attacks.
Storage in the EVM is linked to the concept of gas, as every operation involving storage – whether it is writing new data or modifying existing data – incurs a gas cost. This cost is proportional to the storage resources consumed, reflecting the principle that the more network resources (like memory and storage space) a transaction uses, the more it needs to compensate the network. Every smart contract has its own storage space, which is isolated and theoretically unlimited.
The EVM utilizes a key-value store for data storage, with these key-value pairs referred to as storage slots. A slot’s key is the hash of the contract’s address and the hash of its value. A value is a 32-byte (or 256-bit) word, also called an item. Data smaller than 32 bytes can be packed into a single slot, but if it is larger, the transaction will be reverted [14]. Some types of data are stored in multiple slots, such as arrays and mappings which are stored in multiple contiguous slots [15]. Strings are also stored in multiple slots, where one slot stores the length of the string, and the other slots store each 32-byte chunk of the string [13]. A struct, which is a collection of variables, is stored in a single slot if it fits, otherwise, it is stored in multiple contiguous slots. The gas cost of writing to storage increases with the number of slots written to. This means that writing multiple variables, if they fit within a single slot, incurs the same gas cost as writing a single variable.
@@ 3 Related Work
Software patterns offer a promising approach to this challenge, providing well-tested solutions for frequently encountered application development use cases [3]. A family of software patterns focus on blockchain patterns, which include data management, domain-based, or smart-contract related best practices [28]. In the subcategory of smart contract patterns stand upgradability patterns, which focus on a set of recurring software design approaches to maintain decentralized applications when adding new features or fixing issues are needed [10]. Two recurring patterns emerge through the literature, the Proxy and Diamond patterns [22, 16, 23, 27]
@@
The Proxy pattern enables smart contract updates without changing the contract’s address or requiring data migration[26]. Comprising two contracts – the user-interacted proxy and the logic-holding implementation – the proxy forwards calls to the implementation via delegated calls. Shared storage ensures that if the implementation is updated, the storage persists. To update the implementation, a new contract is deployed, and the proxy is directed to it, eliminating the need for user migration and preventing data loss. Compound, a key DeFi protocol, exemplifies this pattern through multiple upgrades, including Compound III  [^1] https://compound.finance/. Similarly, USDC, a widely-used stable coin, leverages the Proxy pattern for notable upgrades like the transition to USDC 2.0, showcasing its adaptability in the evolving digital currency landscape  [^2] https://www.coinbase.com/blog/usdc-v2-upgrading-a-multi-billion-dollar-erc-20-token.
Figure 1: Diamond Pattern Diagram
@@
The Diamond pattern, depicted in Figure 1 is a more upgradeable version of the Proxy pattern. This pattern solves the maximum contract size problem which is a limitation of the Proxy pattern. Indeed, logic can be separated into small contracts referred to as facets. A main contract, referred to as the implementation or diamond contract, points at the different facets to retrieve the applicable logic. The diamond contract can be upgraded by adding, replacing or removing facets. The Diamond pattern is being utilized by a diverse array of projects, as documented in Nick Mudge’s Awesome Diamonds repository[25]. For example, Aavegotchi, a Non-Fungible Token (NFT) based gaming protocol, employs a single Diamond pattern with eight distinct facets  [^3] https://www.aavegotchi.com/. Each facet serves a specific purpose, contributing to the protocol’s overall functionality. This implementation illustrates how the Diamond pattern can efficiently organize and manage complex functionalities within a single contract structure, an essential feature for sophisticated blockchain applications like gaming protocols.
Two studies focus on gas cost efficiency strategies in smart contracts development. Zarir et al.’s work [33] focuses on transaction parameters rather than architectural design, while the study by Di et al. [12] identifies smart contract coding metrics impacting gas costs but do not specifically focus on upgradeable smart contracts, which possess unique functionalities like proxy pointers and proxy contract management.
@@
A set of studies provide insight on upgradeable smart contract patterns. Kannengiesser et al. [20] conduct a study on key smart contract development challenges across various distributed ledger technology (DLT) protocols, including Ethereum, Hyperledger, and EOSIO. They highlight upgradeability as a significant challenge and reference two upgradeable smart contract patterns, namely the Diamond (referred to as the Façade pattern) pattern and the Proxy pattern. However, the paper lacks an in-depth analysis of these patterns. Two papers [22, 16] identify the Proxy pattern but do not detail the forces, advantages, drawbacks, or gas costs considerations of this pattern. There is no mention of the Diamond pattern. Two other works present a set of upgradeable smart contract patterns, including the Proxy and Diamond patterns [23, 27]. However, the papers do not provide an in-depth analysis of the functional and non-functional properties of these patterns, nor gas costs behaviors.
In summary, a research gap is identified regarding a thorough study of the Proxy and Diamond patterns especially in terms of a gas costs analysis.
@@ 4.1 Protocol
The initial smart contract deployment can be one of the most expensive in terms of gas costs in a smart contract life cycle. Therefore, assessing the deployment costs associated with various patterns is essential. Additionally, upgrades often necessitate deploying additional smart contracts, making it crucial to compare the deployment costs of each subsequent upgrade.
To do so, we propose to develop a file notarization scenario which is a standard use of smart contracts to ensure the integrity of critical data such as diploma or scientific workflows [6, 11, 21]. The scenario will be implemented using the Proxy, and Diamond patterns, as well as a reference monolithic non-upgradeable smart contract referred to as the Classic pattern.
For each pattern, the initial deployment of the notarization smart contract is referred to as Version 1. We will then proceed to two sequential upgrades of the application, a minor one (Version 2) and a major one (Version 3), to evaluate the gas costs. Features will be added as updates are made. Figure 2 demonstrates the upgrades for each pattern.
Figure 2: Evolution of the Notarization Application Across Versions by Patterns
Then the analysis turns to comparing the average gas cost of each function, for seeing which pattern is the less gas intensive when it comes to execution. Each function is accompanied by a test case that is run hundreds of times, to simulate real-world usage of this file notarization application over an extended period and to check the robustness of the code.
@@ 4.2 Implementation
The code will be written in Solidity, the main language used to develop decentralized application on EVM blockchains. For the experiment, the Universal Upgradeable Proxy Standard (UUPS) Proxy pattern is used because it is the recommended standard at the time of writing by OpenZeppelin [7]. For this study, the Diamond pattern implemented by SolidState is used. It utilizes Nick Mudge’s gas-efficient Diamond 2 model [29, 24].
4.2.1 Version 1: Basic Notarization Application
In the first version of the notarization use case, files are represented by a mapping between their name and their hash. The logic will contain the following functions.
  • addFile — A function that notarizes a file on the blockchain. It will modify the state of the contract.
  • getFileName — A view function that returns the name of the file.
  • getFileHash — A view function that returns the hash of the file.
  • compareHash — A pure function that compares two hashes passed as parameters. This operation is ideally performed off-chain, but it will be used here to demonstrate the gas cost of a pure function.
4.2.2 Version 2: Updatable File
@@
In case of file modifications, a function for updating it on the blockchain is needed. The second version of this notarization application will contain the updateFile function, added on top of the previous version. This function updates the file hash of a notarized file and so modifies the state of the contract. This version can be seen as a minor update.
4.2.3 Version 3: Access Control
@@
For security reasons, access control is mandatory, so that only the owner of a file can modify or delete it. This involves creating a File structure, containing the owner’s address, the hash of the file’s contents, the creation timestamp and the last modification timestamp. The previous mapping will be replaced by one between the file’s name and its File structure instance. Then, the addFile and updateFile functions are modified in order to work with this new File structure, and access control is added, where the new logic smart contract ensures that the caller interacts only with his own files. This version can be seen as a major update.
Unit tests are written using Foundry, a smart contract development framework built in Rust [19]. This toolkit is chosen because it compiles and executes faster than other well-known frameworks such as Truffle and Hardhat. Moreover, it eliminates the need for an additional programming language for writing tests and scripts, unlike Truffle and Hardhat with JavaScript, and generates easily gas reports per smart contract function.
The addFile function, in particular, has three test cases with different configurations for the file name parameter to ensure comprehensive coverage. As mentioned earlier, these test cases are conducted repeatedly, with each iteration in the three different configurations of the addFile file name parameter following these criteria: (i) The file name length is incremented by one character, (ii) the file name length remains constant, but the last character varies, (iii) both the file name and hash are unchanged. Foundry autonomously simulates an EVM blockchain environment during test execution. It deploys the contracts and carries out the testing scenarios within this emulated setting. The testing utilizes the default configuration of this local blockchain.
The entire source code, tests and results here are available in the accompanying source code repository. [^4] https://github.com/opixelum/proxy-diamond-patterns-gas-analysis

### v2

# A Comparative Gas Cost Analysis of Proxy and Diamond Patterns in EVM Blockchains for Trusted Smart Contract Engineering

@@ Abstract
Blockchain applications are witnessing rapid evolution, necessitating the integration of upgradeable smart contracts. Software patterns have been proposed to summarize upgradeable smart contract best practices. However, research is missing on the comparison of these upgradeable smart contract patterns, especially regarding gas costs related to deployment and execution. This study aims to provide an in-depth analysis of gas costs associated with two prevalent upgradeable smart contract patterns: the proxy and diamond patterns. The proxy pattern utilizes a proxy pointing to a logic contract, while the diamond pattern enables a proxy to point to multiple logic contracts. A comparative analysis of gas costs for both patterns is conducted and compared to a traditional non-upgradeable smart contract. From this analysis, a theoretical contribution is derived in the form of two consolidated blockchain patterns and a corresponding decision model.
@@ 1 Introduction
Smart contracts are pivotal for orchestrating digital transactions in a reliable and secure manner in blockchain platforms [19]. Smart contracts are particularly important on Ethereum and similar blockchain platforms, where they are used in diverse areas including digital finance or industrial traceability.
@@
As blockchain applications evolve, the need for smart contracts to be not only secure but also upgradeable becomes apparent [8]. In this context, a set of upgradeable smart contract patterns, including the proxy and diamond patterns, have surfaced to answer the lack of classical smart contracts’ adaptability whose logic cannot be changed once deployed [22, 27]. The proxy pattern uses a proxy contract to delegate calls to an implementation contract, providing a flexible and upgradeable solution. The diamond pattern, introduced in EIP-2535, addresses concerns like contract size limitations and facilitates enhanced maintainability and versioning through multiple implementation contracts.
@@
Research has focused on identifying upgradeable smart contracts pattern families, pointing towards proxy and diamond-based strategies  [21, 14, 22, 27]. However, the papers do not provide an in-depth analysis of the functional and non-functional properties of these patterns, nor gas costs behaviors. Hence, a research gap is identified regarding a thorough study of the proxy and diamond patterns, especially in terms of a gas costs analysis.
@@
To address this issue, this paper aims to answer the following research questions: (RQ1) How do the classic, proxy, and diamond patterns differ in terms of gas consumption, scalability, and ease of use? And (RQ2) What implications do these differences have for the development of blockchain applications, considering the traditional classic pattern as a baseline?
@@
In this paper, we contribute to the literature through a unified approach to compare gas costs in upgradeable smart contracts. We leverage this methodology to provide a comparative gas cost analysis of the proxy and diamond patterns in EVM blockchains, compared against a monolithic non-upgradeable smart contract, which is used as a baseline. Based on these results, we derive a theoretical contribution in the form of two smart contract patterns adhering to the Alexandrian form format following the standard proposed by Christopher Alexander [1]. These patterns include the results of our comparative analysis and contribute to the broader understanding of upgradeable smart contract patterns and their use in blockchain applications. Based on these patterns, a decision model for using these patterns is proposed, emphasizing functional and non-functional properties for each design decision.
@@
The remainder of this paper is structured as follows. Section 2 introduces key concepts related to smart contracts and blockchain patterns, and section 3 presents studies already made on linked concepts. Section 4 presents the method used to carry a comparative analysis on proxy versus diamond gas costs. Section 5 presents the results of the tests made on the different patterns. Section 6 leverages these findings to propose two consolidated proxy and diamond patterns as well as a decision model for using these patterns. Section 7 finally concludes the paper with a summary and a discussion of the results and some considerations for future work.
@@ 2.1 Smart Contracts
A smart contract is a program hosted on a blockchain network [30]. When a smart contract executes, its updated state (or storage) is registered into a transaction. Then, that transaction is stored in the blockchain ledger making it immutable and tamper-proof [4]. More precisely, the final validated results are stored in a Merkle Patricia trie whose nodes correspond to an account or a smart contract. A smart contract comprises both variables and functions. Smart contract functions can execute arbitrary code, access the state of the variables and optionally update them. The default function’s mutability gives the right to read and modify state variables. However, function mutability can be constrained to add more security when it comes to accessing these variables. On the one hand, a pure function cannot read nor modify the state of the contract; it is only used to compute a value, often using the parameters passed to it. On the other hand, a view function can read the state of the contract, but cannot modify it. Smart contracts are immutable, meaning that once deployed, they cannot be modified [31, 7]. This is a security feature, as it prevents malicious actors from modifying the code, so that the contract can be trustable and unbreakable. However, this also means that if a bug is found in the code, it cannot be fixed, and the only solution is to deploy a new contract. Also, if a new feature is added to the contract, the only way to do it is to deploy a new contract. This is a problem for users, as they have to migrate to the new contract, and for developers, as they have to maintain multiple contracts. To address this issue, several patterns have emerged, which are discussed later in the paper [8, 22].
@@ 2.2 Gas and Storage in the Ethereum Virtual Machine (EVM)
The EVM is a Turing-complete virtual machine that runs on every node of the Ethereum network and other EVM-based blockchains. It provides a secure and isolated environment for smart contract execution [16]. Gas is the unit of computation in the EVM, and every operation performed by the EVM has a gas cost associated with it [15]. The user pays this cost in the form of gas fees, to the miner/validator who executes the smart contract. It is the miner’s incentive to execute the smart contract and record the results on the blockchain, but also a way to prevent spam and denial of service attacks. Storage in the EVM is linked to the concept of gas, as every operation involving storage – whether it is writing new data or modifying existing data – incurs a gas cost. This cost is proportional to the storage resources consumed, reflecting the principle that the more network resources (like memory and storage space) a transaction uses, the more it needs to compensate the network. Every smart contract has its own storage space, which is isolated and theoretically unlimited.
Smart contracts’ storage utilizes a key-value store, with these key-value pairs referred to as storage slots. A key for a storage slot is determined by the index of the slot, which is numbered contiguously from 0 to $12^{256}-$ 1. A value is a 32-byte (or 256-bit) word, also called an item. Data smaller than 32 bytes can be packed into a single slot, but if it is larger, the transaction is reverted [12]. Some types of data are stored in multiple slots, such as arrays and mappings which are stored in multiple contiguous slots [13]. Strings are also stored in multiple slots, where one slot stores the length of the string, and the other slots store each 32-byte chunk of the string [11]. A struct, which is a collection of variables, is stored in a single slot if it fits, otherwise, it is stored in multiple contiguous slots. The gas cost of writing to storage increases with the number of slots written to. This means that writing multiple variables, if they fit within a single slot, incurs the same gas cost as writing a single variable.
@@ 3 Related Work
(a) Proxy pattern diagram.
(b) Diamond pattern diagram.
Figure 1: Illustration of the Proxy and Diamond patterns
Software engineering patterns provide well-tested solutions for frequently encountered application development use cases [2]. Standardized formatting such as the Alexandrian Form Format systematically include for each pattern the description, the forces or tradeoffs at stake, the benefits and drawbacks, and their main applications. A family of patterns focusing on blockchain patterns has recently emerged in the literature [33, 28]. The later include smart-contract related best practices such as upgradability patterns which refer to decentralized application maintenance strategies that can be applied for adding or updating features [8]. Two recurring upgradability patterns are identified in the literature, namely the proxy and diamond patterns [21, 14, 22, 27].
@@
The proxy pattern, pictured in Figure 1(a) enables smart contract updates without changing the contract’s address or requiring data migration[25]. It consists of two contracts, the user-interacted proxy and the logic-holding implementation. The proxy forwards calls to the implementation via delegated calls. Shared storage ensures that if the implementation is updated, the storage persists. To update the implementation, a new contract is deployed, and the proxy is directed to it, eliminating the need for user migration and preventing data loss. Compound, a key DeFi protocol, exemplifies this pattern through multiple upgrades, including Compound III  [^1] https://compound.finance/.
@@
The diamond pattern, depicted in Figure 1(b) is a more upgradeable version of the proxy pattern. This pattern solves the maximum contract size problem, which is a limitation of the proxy pattern. Indeed, logic can be separated into small contracts referred to as facets. A main contract, referred to as the implementation or diamond contract, points at the different facets to retrieve the applicable logic. The diamond contract can be upgraded by adding, replacing or removing facets. The diamond pattern is used in a diverse array of projects, as documented in Nick Mudge’s Awesome Diamonds repository[24]. For example, Aavegotchi, a Non-Fungible Token (NFT) based gaming protocol, employs a single diamond pattern with eight distinct facets  [^2] https://www.aavegotchi.com/.
@@
A set of studies provide insight on upgradeable smart contract patterns. Kannengiesser et al. conduct a study on key smart contract development challenges across various distributed ledger technology (DLT) protocols, including Ethereum, Hyperledger, and EOSIO [19]. They highlight upgradability as a significant challenge and reference two upgradeable smart contract patterns, namely the diamond (referred to as the façade pattern) pattern and the proxy pattern. However, the paper lacks an in-depth analysis of these patterns. Two papers identify the proxy pattern but do not detail the forces, advantages, drawbacks, or gas costs considerations of this pattern [21, 14]. There is no mention of the diamond pattern. Two other works present a set of upgradeable smart contract patterns, including the proxy and diamond patterns [22, 27]. However, the papers do not provide an in-depth analysis of the functional and non-functional properties of these patterns, nor gas costs behaviors. Additionally, it is to note that two studies focus on gas cost efficiency strategies in smart contracts development. Zarir et al.’s work focuses on transaction parameters rather than architectural design [34]. The study by Di et al. identifies smart contract coding metrics impacting gas costs [10]. However, it does not specifically focus on upgradeable smart contracts. These contracts possess unique functionalities like proxy pointers and proxy contract management. In summary, a research gap is identified regarding a thorough study of the proxy and diamond patterns. There is a lack of studies about gas costs analysis and formalization of the proxy and diamond patterns, especially using the Alexandrian form format.
@@ 4.1 Protocol
Figure 2: Evolution of the notarization application across versions by patterns.
Smart contract deployments are one of the most expensive operations in terms of gas costs. Therefore, assessing the deployment costs associated with various patterns is essential. Additionally, upgrades often necessitate deploying additional smart contracts, making it crucial to compare the deployment costs of each subsequent upgrade. The protocol leverages a file notarization scenario, a standard use of smart contracts to ensure the integrity of critical data such as diplomas or scientific workflows [5, 9, 20]. The scenario is implemented using the proxy and diamond patterns, as well as a reference monolithic non-upgradeable smart contract referred to as the classic pattern. For each pattern, the initial deployment of the notarization smart contract is referred to as version 1. The study then proceeds to two sequential upgrades of the application, a minor one (version 2) and a major one (version 3), to evaluate the gas costs. Features are added as updates are made. Figure 2 demonstrates the upgrades for each pattern. Then the analysis turns to comparing the average gas cost of each function, for seeing which pattern is the less gas intensive when it comes to execution. Each function is accompanied by a test case ran hundreds of times, to simulate real-world usage of this file notarization application over an extended period and to check the robustness of the code.
@@ 4.2 Implementation
The code is written in Solidity, the main language used to develop decentralized application on EVM blockchains. For the experiment, the Universal Upgradeable Proxy Standard (UUPS) proxy pattern is used because it is the recommended standard at the time of writing by OpenZeppelin [6]. For this study, the diamond pattern implemented by SolidState is used. It utilizes Nick Mudge’s gas-efficient Diamond 2 model [29, 23] in a plug-and-play fashion: the developer only needs to import the diamond made by SolidState without any configuration needed.
Version 1: Basic Notarization Application. In the first version of the notarization use case, files are represented by a mapping between their name and their hash. The logic encompasses several functions. The function addFile is designed to notarize a file on the blockchain and modifies the state of the contract. The getFileName function is a view function that returns the name of the file. Similarly, getFileHash is a view function that provides the hash of the file. Lastly, compareHash is a pure function for comparing two hashes passed as parameters; while this operation is ideally performed off-chain, it is included here to demonstrate the gas cost of a pure function.
@@
Version 2: Updatable File. In case of file modifications, a function for updating it on the blockchain is needed. The second version of this notarization application contains the updateFile function, added on top of the previous version. This function updates the file hash of a notarized file and so modifies the state of the contract. This version can be seen as a minor update.
@@
Version 3: Access Control. For security reasons, access control is mandatory, so that only the owner of a file can modify or delete it. This involves creating a File structure containing the owner’s address, the hash of the file’s contents, the creation timestamp and the last modification timestamp. The previous mapping is replaced by one between the file’s name and its File structure instance. Then, the addFile and updateFile functions are modified to work with this new File structure, and access control is added, where the new logic smart contract ensures that the caller interacts only with his own files. This version can be seen as a major update.
4.3 Unit Tests
As mentioned in section 4.1, each function has a unit test that is run hundreds of times (704 iterations for the addFile function across all unit tests of version 3, for example). It calls the function with random parameters (file name, hash, etc.) and checks that the result is correct. This is done to simulate real-world usage of this file notarization application. An expected result is computed before the function call, and the actual result is compared to it. In the addFile unit test, we expect the smart contract to add the file name and hash to its storage, using either the proxy or diamond pattern. The updateFile unit test should update the file hash in the contract’s storage, while the deleteFile test ensures the file’s removal. In the compareHash test, the outcome should be true for matching hashes and false otherwise. The getFileName, getFileHash, getFileOwner, getFileCreatedAt, and getFileLastModifiedAt unit tests respectively verify the return of the file’s name, hash, owner address, creation timestamp, and last modification timestamp. Finally, the getFileDetails test checks for the return of all the previously mentioned file properties.
Testing each single function is important to ensure that the code is robust and that it does not break when upgrading the smart contract. These tests are also conducted to compare the gas costs across various patterns, specifically examining each type of function, including pure, view, and state-modifying functions. By testing pure functions, the gas costs of computations without storage access can be assessed and compared across patterns. View functions are tested to evaluate the gas costs of computations with storage access. Finally, state-modifying functions are tested to assess the gas costs of storage writes.
Smart contracts used for this study are developed using Foundry [17]. Foundry plays a critical role in this testing process by autonomously simulating an Ethereum Virtual Machine (EVM) blockchain environment. During test execution, it deploys the contracts and carries out the testing scenarios within this emulated setting, utilizing the default configuration of this local blockchain. The entire source code, tests and results here are available in the accompanying source code repository  [^3] https://anonymous.4open.science/r/proxy-diamond-patterns-gas-analysis.